OpenStudy (anonymous):

hey its me again which of the following rational numbers are equal to 3/5?: 12/20, -3/-5, 0.3 or 6/10

jimthompson5910 (jim_thompson5910):

3/5 = 12/20 3/5 = -3/-5 3/5 = 6/10

jimthompson5910 (jim_thompson5910):

A quick check with all of them is plugging them into a calculator to see that they all equal 0.6

OpenStudy (anonymous):

\[\left\{\frac{12}{20},\frac{-3}{-5},0.3,\frac{6}{10}\right\}=\{0.6,0.6,0.3,0.6\} \]